OpenStudy (kendricklamar2014):

Simplify: -m^2 * 4r^3 * 12r^-4 * 5m

OpenStudy (kendricklamar2014):

@jigglypuff314 @jim_thompson5910 @Nnesha

jigglypuff314 (jigglypuff314):

know that x^a * x^b = x^(a+b) so like r^3 * r^(-4) = r^(3 + (-4)) = r^( )

OpenStudy (kendricklamar2014):

r = -1

jigglypuff314 (jigglypuff314):

r^(-1) yes

OpenStudy (kendricklamar2014):

So what is the first step to solve the problem?

jigglypuff314 (jigglypuff314):

this is just multiplying variables next is m^2 * m^1 = m^(2+1) = ?

OpenStudy (kendricklamar2014):

= m^3 right?

jigglypuff314 (jigglypuff314):

yes then you just multiply the other numbers (-1) (4) (12) (5)

OpenStudy (kendricklamar2014):

Multiply 1 * 4 * 12 * 5??

jigglypuff314 (jigglypuff314):

negative 1 but yeah :)

OpenStudy (kendricklamar2014):

-1 * 4 * 12 * 5 = -240

jigglypuff314 (jigglypuff314):

good! now combine everything I've had you solve for

OpenStudy (kendricklamar2014):

-240m^3

jigglypuff314 (jigglypuff314):

and the r

OpenStudy (kendricklamar2014):

does it go in the beginning of the answer or the end??

jigglypuff314 (jigglypuff314):

in alphabetical order numbers then letters

OpenStudy (kendricklamar2014):

So like this -240m^3 r^(-1

jigglypuff314 (jigglypuff314):

yes ^_^ I think so

OpenStudy (kendricklamar2014):

Thanks